The showlabels package^{∗}

Norman Gray

                 norman@astro.gla.ac.uk

April 23, 1995

This is the showlabels class option.

This class option introduces no new commands, it just helps you keep track of all the labels by putting the name of the new label into the margin whenever either the \label command is used, or an equation is automatically numbered (by replacing the internal \@eqnnum command). There are therefore no usage instructions. You invoke this package with the command \usepackage{showlabels} in the preamble, and you may give the options [inner] or [outer] (the default) to have the labels placed in the inner or outer margin of the text.

∗This file has version number v1.0, last revised 1994/10/03, documentation dated 1994/10/03